380 pints =how many gallons
Answer by jim_thompson5910(35256)

1 gallon = 8 pints

1/8 gallon = 1 pint

1 pint = 1/8 gallon

380*1 pint = 380*1/8 gallon

380 pints = 47.5 gallons